The equation [ x – 1][x + 1] = a^2 – 2a – 3 can have real solutions for x if a belongs to
 

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
(x^2-1) = a^2-2a-3 \\=>a^2-2a-2 \geq 0 \\=>(a-1)^2-3 \geq 0 \\=>(a-1) \in (-\infty,\sqrt3] \bigcup [\sqrt3,\infty]
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty